OpenStudy (anonymous):

A caterer charges a fixed cost per party, plus an additional charge per guest. If it costs $406 to serve 32 guests and $484 to serve 45 guests, find the fixed cost and the cost per guest. A. $86 cost, $10 per guest B. $156 cost, $8 per guest C. $214 cost, $6 per guest D. $246 cost, $5 per guest

OpenStudy (anonymous):

Equations formed will be: Let x be fixed cost per party and y be cost per guest... x + 32y = 406 x + 45y = 484 Then solve it for x and y..

OpenStudy (anonymous):

A ? (:

OpenStudy (anonymous):

@waterineyes : A ? (:

OpenStudy (anonymous):

No, I think I am wrong in making those equations...

OpenStudy (anonymous):

X + 32Y = 406 X = 406 - 32Y X + 45Y = 484 X = 484 - 45Y X = X, so: 406 - 32Y = 484 - 45Y add 45Y to each side gets you: 406 + 13Y = 484 subtract 406 from each side: 13Y = 78 Y = 6 Put Y = 6 back into the original equations: X + 32(6) = 406 X = 214 X + 45(6) = 484 X = 214 C is the correct answer.

OpenStudy (anonymous):

Answer is C...

OpenStudy (campbell_st):

just look at it this way an addtional 13 guests cost $78.00 or $6 per guest so 406 = f + 32*6 gives the fixed cost of $214 just checking 484 = f + 45*6 f = $214 fixed cost is $214

OpenStudy (anonymous):

Eliminate x through first equation : x = 406 - 32y Now put it in second equation : We get, 13y = 78 So, y = 6 and therefore x = 214.. So, fixed cost is $214 and Cost per Guest is $6...